OpenStudy (anonymous):

Simplify $\left(\frac{103}{301}\right)\left(\frac{103}{301}\right)^{-1} + \left(\frac{605}{506}\right)\left(\frac{605}{506}\right)^{-1}$

OpenStudy (help!!!!):

yeah copy and paste that in the equation button

OpenStudy (anonymous):

\[(103/301) (103/301)^-1+(605/506) (605/506)^-1\]

OpenStudy (radar):

\[$\left(\frac{103}{301}\right)\left(\frac{103}{301}\right)^{-1} + \left(\frac{605}{506}\right)\left(\frac{605}{506}\right)^{-1}$\]

OpenStudy (radar):

1 + 1 = 2 $2

OpenStudy (anonymous):

ok

OpenStudy (help!!!!):

Ha

OpenStudy (anonymous):

now I understand what you mean

OpenStudy (radar):